OpenStudy (anonymous):

Convert the function into vertex form. Show your work. y=x^2+4x+9

OpenStudy (shamil98):

y= a(x-h)^2 + k

OpenStudy (anonymous):

i need work shown i need to understand

OpenStudy (shamil98):

-b/2a = the axis of symmetry and x coordinate to your vertex

OpenStudy (shamil98):

-4/2 = -2

OpenStudy (shamil98):

Then plug -2 into your equation to find the y coordinate.

OpenStudy (shamil98):

y = -2^2 + 4(-2) + 9 y = 4 - 8 + 9 y = -4 + 9 y = 5

OpenStudy (***[isuru]***):

convert x^2+4x+9 in to a complete square...